OpenStudy (anonymous):

21. (a + 2b + 3c) − (4a + 6b − 5c) is equivalent to: A. −4a − 8b − 2c B. −4a − 4b + 8c C. −3a + 8b − 2c D. −3a − 4b − 2c E. −3a − 4b + 8c

OpenStudy (anonymous):

I got D.

OpenStudy (anonymous):

you are right

OpenStudy (michele_laino):

hint: \[\begin{gathered} \left( {a + 2b + 3c} \right) - \left( {4a + 6b - 5c} \right) = \hfill \\ = a + 2b + 3c - 4a - 6b + 5c \hfill \\ \end{gathered} \]